OpenStudy (anonymous):

(v-4/v+8)+(v+8/v-12)

OpenStudy (anonymous):

\[ \Large {\left( v - \dfrac {4}{v} + 8 \right) + \left( v + \dfrac {8}{v} - 12 \right) = v + v - \dfrac {4}{v} + \dfrac {8}{v} + 8 - 12 = } \]\[ \Large{\boxed {2v + \dfrac {4}{v} - 4}}. \]